OpenStudy (anonymous):

Which equation is related to the equation shown? 5x = 35 A. x = 35 + 5 B. x = 35 ÷ 5 C. x = 35 − 5 D. x = 35 • 5

OpenStudy (midhun.madhu1987):

Divide both sides by 5 and you get the answer..

OpenStudy (anonymous):

The answer is B. 5x = 35, divide both sides by 5 and you get 5x/5 = 35/5, which in return is x = 35/5

OpenStudy (anonymous):

Try B, if your equation is 5x=35 then you need to divide 35 by 5. Then go through your options and find the same sum as you got when you divided and that's the correct one. Hope I helped you understand it better.